OpenStudy (anonymous):

Solve. x + 5 – 8x = 26 (Points : 4) 5 3 –3 –5

OpenStudy (anonymous):

x + 5 - 8x = 26 First, you want to combine like terms. (x - 8x) + 5 = 26 -7x + 5 = 26 Now, to isolate the variable on one side, just subtract 5 from both sides. -7x + 5 -5 = 26 -5 -7x = 21 Now, just divide each side by -7. -7x/-7 = 21/-7 x = ?

OpenStudy (anonymous):

Beat me to it lol

OpenStudy (anonymous):

Bahaha, sorry. :P I gave you a medal. (:

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

You're welcome!